A circular flower garden has an area of 314 m2 . A sprinkler at the centre of the garden can cover an area that has a radius of 12 m. Will the sprinkler water the entire garden? (Take π = 3.14)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,
Area of the circular flower garden = 314 m^2
Sprinkler at the centre of the garden can cover an area that has a radius = 12 m
Area of the circular flower garden =πr^2
314 = 3.14 × r^2
314/3.14 = r^2
r^2= 100
r =√100
r = 10 m
∴Radius of the circular flower garden is 10 m.
Since, the sprinkler can cover an area of radius 12 m
Hence, the sprinkler will water the whole garden.

Thanks